geerky42 (geerky42):

How many perfect squares divide the number 4!·5!·6! ?

geerky42 (geerky42):

How did you get that answer?

OpenStudy (anonymous):

take that back hold on.....what else did it say

geerky42 (geerky42):

Nothing else.

OpenStudy (anonymous):

is it like 4*5*6=120

geerky42 (geerky42):

120 is a reasonable answer, but your solution doesn't make sense. It is asking for number of perfect squares that 4!·5!·6! are divisible of.

OpenStudy (anonymous):

You'll want to prime factor the number first:\[4!\cdot 5!\cdot 6!=(4!)^3\cdot 5^2\cdot 6=24^3\cdot 5^2\cdot 6=(3\cdot 2^3)^3\cdot 5^2\cdot (2\cdot 3)\]\[=2^{10}\cdot 3^4\cdot 5^2\]

OpenStudy (anonymous):

Now start counting the ways you can factor squares out. Its a little tedious (unless there is a clever counting argument available, im not good with combinatorics). Just looking at the powers of 2, its clear that:\[2^2,2^4,2^6,2^8,2^{10}\]are all squares that divide your numbers.

OpenStudy (anonymous):

Maybe this will work, factor the primes as such:\[(2^2)^5\cdot (3^2)^2\cdot (5^2)\]Then its clear that the number of square divisors is:\[(5+1)\cdot (2+1)\cdot (1+1)=36\]

geerky42 (geerky42):

Smart. Thanks!